OpenStudy (anonymous):

Simplify radical expression.. square root 10/81

OpenStudy (apoorvk):

\[\sqrt{\frac {10}{81}} =\sqrt{\frac {2 \times 5}{3 \times 3}} = \frac 1 3 \sqrt{2 \times 5} = \frac{\sqrt{10}}3\]

OpenStudy (anonymous):

thanks! what about this one \[\sqrt{6} + 2 \sqrt{6}\]

Parth (parthkohli):

@apoorvk \(\sqrt{81} \ne 3\)

OpenStudy (apoorvk):

Omg thanks Parth, I realised it yeah... Sorry @zebra11, a bit of a correction, \(\sqrt{81} = \sqrt{9 \times 9} = 9\) can you apply this correction by your yourself?

OpenStudy (anonymous):

Yeah I got it.. what about the \[\sqrt{6} + 2 \sqrt {6}\]

OpenStudy (apoorvk):

hmm, \(\sqrt6\) is common to both terms, so if you take it out as common, you'd get: \((1+2)\sqrt6 = 3\sqrt6 \) helps?

OpenStudy (anonymous):

yes. thanks!

OpenStudy (apoorvk):

no worries..